# Quantum Advantage in Distributed Sensing with Noisy Quantum Networks

### v2

@@ Abstract
We show that quantum advantage in distributed sensing can be achieved with noisy quantum networks. When using depolarized GHZ states as the probe, we derive a closed-form fidelity threshold to achieve advantage over the optimal local sensing strategy. The threshold indicates that while entanglement is needed for this quantum advantage, genuine multipartite entanglement is generally unnecessary. We further explore the impacts from imperfect local entanglement generation and local measurement constraint, and our results imply that the quantum advantage is more robust against quantum network imperfections than local operation errors. Finally, we demonstrate that the quantum advantage in distributed sensing can be achieved with a three-node quantum network using practical protocol stacks through simulations with SeQUeNCe, an open-source, customizable quantum network simulator.
@@
Introduction.—Distributed quantum sensing (DQS)[1, 2, 3] is one of the most important applications of quantum networks [4, 5]. It is expected to surpass classical sensing techniques in areas ranging from magnetometry [6, 7, 8, 9], phase imaging [10], precision clocks [11], energy applications [12], all the way to the exploration of fundamental physics [13, 14], including search for dark matter [15] and measuring stability of fundamental constants [16].
Over the past decade, DQS has attracted great efforts on the derivation of its ultimate limit [17, 18, 19, 20, 21, 22]. Various DQS protocols have also been proposed under ideal conditions [23, 24, 25, 26]. On the other hand, as the principle of DQS has been experimentally demonstrated in small-scale matter-based quantum systems [27], and the experimental implementation of entanglement distribution networks is quickly advancing [28, 29, 30], it is crucial to analyze the feasibility of DQS with realistic, noisy quantum networks. Despite its importance and necessity, the analysis of DQS with imperfection, especially the state preparation error that is unavoidable in realistic quantum networks, has only emerged in recent years and is insufficient. For instance, Sekatski et al. [31] and some later works [32, 33, 34] considered DQS with noise in signals, rather than the unavoidable state preparation error over a realistic quantum network. Fadel et al. [35] studied partitioning atomic ensembles for DQS, but again did not consider state preparation noise, and the assumed physical scenario is still small-scale. Van Milligen et al. [36] considered a quantum network scenario with focus on the probabilistic nature of network operations.
In this work, we focus on practical DQS scenarios and study the possibility of realizing quantum advantage in DQS with noisy quantum networks. We first specify the initial probe state, the parameter encoding process, and the estimation objective. We then derive closed-form upper bounds for DQS estimation accuracy and theoretically demonstrate the possibility of quantum advantage over the optimal local sensing protocol. The impacts from imperfect local entanglement generation when scaling up the number of local sensors, and local measurement constraint are also analyzed. In addition, we perform a quantum network simulation which shows that DQS quantum advantage can in principle be achieved by realistic quantum network configurations.
@@
Under the aforementioned unitary encoding channel, the optimal probe state for this problem has been shown to be the global GHZ state  [2, 1]. Therefore, we choose this global GHZ state as the target initial probe state to prepare in the quantum network of sensors. The preparation of the probe state over a quantum network can be decomposed into two steps. Firstly the quantum network will distribute a $d$-qubit GHZ state across all $d$ sensor nodes, with each node having one qubit. Each node can then perform local entanglement generation [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48] to extend the size of the global GHZ state: By entangling $(n-1)$ additional quantum sensors per node, the global GHZ state will have $nd$ qubits in total. Given noisy quantum networks, we consider that the prepared initial probe state is mixed. Motivated by Pauli twirling [49, 50, 51] which is able to transform error channels into Pauli channels, and stabilizer twirling [52] which can cancel off-diagonal density matrix elements under a stabilizer basis, we assume that the initial probe state is GHZ-diagonal: $\rho_{0}=\sum_{a}\lambda_{a}|\psi_{a0}\rangle\langle\psi_{a0}|$ with all $\lambda_{a}$ being non-zero, where $a$ is the index of the pure GHZ basis state $|\psi_{a0}\rangle$. This Pauli error model has also been utilized in a recent study of single-parameter quantum metrology with graph states [53]. According to the assumed GHZ-diagonal form of initial probe state, we can analytically derive the quantum Fisher information (QFI) [54] which characterizes the lower bound of parameter estimation variance, the quantum Cramér-Rao bound (QCRB).
Quantum Fisher information.—To derive the QFI for the average of all local parameters, we start with the QFI matrix (QFIM) [55, 56, 57, 58] for the $d$ local parameters. According to the assumed parametrization unitary, for a general GHZ-diagonal $nd$-qubit initial probe state $\rho_{0}$ where each of the $d$ sensor nodes holds $n$ qubit sensors, the QFIM is always proportional to a matrix of ones:
$\mathcal{F}(x)=(1-C)n^{2}\begin{pmatrix}1&1&\dots&1\\ 1&1&\dots&1\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\dots&1\\ \end{pmatrix},$ (1)
where $\mathcal{F}(x)$ denotes the QFIM for parameters $x$, and $C$ captures the quality of the initial probe state, and it is analytically calculated as:
@@
Our parameter to estimate is $\theta_{1}=v_{1}^{T}x$ where $v\propto(1,\dots,1)^{T}$. For concreteness, we follow the convention in [2, 1] and choose $v_{1}=(1,\dots,1)/\sqrt{d}$ which is normalized under the 2-norm. Then we are able to transform the QFIM for the new global parameter $\theta_{1}$, where the transformation could be operationally realized by constructing additional $(d-1)$ normalized vectors $v_{2},\dots,v_{d}$, s.t. $v_{i}^{T}v_{j}=\delta_{ij}$ [55]. Consequently, we have:
@@
where $\theta=(v_{1}^{T}x,\dots,v_{d}^{T}x)$, and the only non-zero entry is the QFI for $\theta_{1}$ of our interest: $\mathcal{F}(\theta_{1})=d(1-C)n^{2}$. The detailed derivation for the above results can be found in Sec. II of the Supplemental Material [59].
@@
It is well known that the GHZ state is the optimal probe state for local phase estimation [60], and indeed using separate GHZ states on each sensor node is the best local strategy (without any quantum communication between sensor nodes) for estimating the average of local parameters [2, 1]. The variance of this indirect estimation can be calculated through propagation of error [57]: $\mathrm{Var}_{\mathrm{local}}(\hat{\theta}_{1})=\sum_{l=1}^{d}\left(\frac{\partial\theta_{1}}{\partial x_{l}}\right)^{2}\mathrm{Var}_{\mathrm{local}}(\hat{x}_{l})$, where we use the QCRB of local GHZ state with $n$ qubits so that $\mathrm{Var}_{\mathrm{local}}(\hat{x}_{l})=1/(n^{2}\mu)$, where $\mu$ is the number of samples. Noticing that $\partial\theta_{1}/\partial x_{l}=1/\sqrt{d}$, we immediately have $\mathrm{Var}_{\mathrm{local}}(\hat{\theta}_{1})=1/(n^{2}\mu)$. Therefore, a quantum advantage of entangling sensor nodes with quantum networks is achievable when:
$\eta=d(1-C)>1.$ (4)
Next, we will evaluate $\eta$ for $nd$-qubit GHZ state created from the initial noisy $d$-qubit GHZ state with noiseless and noisy local entanglement generation, respectively.
Noiseless local entanglement generation.—We first consider noiseless local entanglement generation, which can be interpreted as applying perfect CNOT gates between the qubit sensor initially entangled with other sensor nodes and other local qubit sensors prepared in $|0\rangle$ state to entangle, with the former being the control. This assumption allows us to further decouple the network imperfections and local errors, and thus to understand the limit on the amount of network imperfections beyond which there cannot be any quantum advantage for DQS. Without loss of generality, we assume that the initial noisy $d$-qubit GHZ state is full-rank, and in the form of a depolarized GHZ state, i.e. a mixture of pure GHZ state and maximally mixed state, $\rho_{\mathrm{dp}}(F)$, where $F$ is the fidelity to the pure GHZ state. We can then derive a closed-form fidelity threshold, below which there will be no quantum advantage in DQS:
@@
We see that in the asymptotic regime of large $d$, the above fidelity threshold reduces to $1/d$. Moreover, the quantum advantage in DQS has a close relationship with quantum entanglement, and the entanglement properties of the depolarized GHZ state have been well studied: $d$-qubit depolarized GHZ states are not completely separable if $F>3/(2^{d}+2)$ [61], and genuine multipartite entangled (GME) if $F>1/2$ [62]. It can be shown that $F_{\mathrm{th,dp}}>3/(2^{d}+2)$ for $d\geq 2$, which means that entanglement is necessary for DQS quantum advantage. Meanwhile, except for the special case of $\left.F_{\mathrm{th,dp}}\right|_{d=3}\approx 0.50963>1/2$, $F_{\mathrm{th,dp}}<1/2$ for $d>3$, which suggests that GME is in principle unnecessary to demonstrate quantum advantage in DQS. Further details can be found in Sec. III.A of the Supplemental Material [59].
@@
Figure 1: Visualization of $\eta_{\mathrm{dp}}$ as function of the number of local qubits $n$, for some choices of initial fidelity $F$ and local entanglement generation quality $k$. We fix $d=3$ while varying $F$ and $k$. Dashed curves correspond to $F=0.8$; solid curves correspond to $F=0.9$; dot-dashed curves correspond to $F=0.99$. Blue color represents $k=0.9999$; yellow color represents $k=0.999$; green color represents $k=0.99$.
@@
Imperfect local entanglement generation.—We now start to include imperfections in local entanglement generation. Specifically, we consider the following phenomenological model to reflect the fidelity decrease when adding more qubits to the GHZ state: We assume that the state after local entanglement generation is an $nd$-qubit depolarized GHZ state, while the fidelity is modified according to the number of local qubits as $F(n)=k^{n-1}F$, where $F$ is the fidelity of the initial $d$-qubit GHZ state, and $k\in(0,1)$ is a parameter which describes the quality of local entanglement generation and thus the larger the better. The intuition behind this phenomenological model is the usage of noisy CNOT gates to generate GHZ states, and $k$ can be interpreted as gate fidelity. We can then express $C$ in Eqn. 2 as:
It can be shown that, as intuitively expected, $C_{\mathrm{dp}}$ decreases monotonically as initial fidelity $F$, number of local parameters $d$, and local entanglement generation quality $k$, increase for all $n\geq 1$, if $F,k>2^{-d}$. On the other hand, the dependence of $C_{\mathrm{dp}}$ on number of local quantum sensors $n$ is more complicated. We visualize $\eta_{\mathrm{dp}}=d(1-C_{\mathrm{dp}})$ with varying $n$ under different parameter choices in Fig. 1.
@@
From the plot we can verify that the relative advantage $\eta_{\mathrm{dp}}$ is larger when $F,k$ increase. We can also see non-monotonic behavior of $\eta_{\mathrm{dp}}$ when $k$ is sufficiently high in comparison to $F$, which indicates an increase in relative advantage for small local sensor numbers. The intersection point at which $\eta_{\mathrm{dp}}=1$ determines the maximal number of local quantum sensors $n_{\mathrm{max}}$ for quantum advantage over the optimal local sensing strategy to be potentially demonstrated. We observe that $n_{\mathrm{max}}$ does not change much when varying $F$ given fixed $k$, but it changes significantly when fixing $F$ and varying $k$, which suggests that the initial fidelity $F$ has less impact on $n_{\mathrm{max}}$ than the local entanglement generation quality $k$. This can be understood by considering the fidelity threshold for the $nd$-qubit depolarized GHZ state to be advantageous over the optimal local strategy, which also scales as $1/d$ and is almost independent of $n$. Imperfect local entanglement generation will decrease the GHZ fidelity to the threshold at $n_{\mathrm{max}}$. Therefore, we expect $k^{n_{\mathrm{max}}-1}F\approx 1/d$, which gives $n_{\mathrm{max}}\approx-\ln(dF)/\ln(k)$. This also implies robustness of the quantum advantage in DQS against the imperfection of probe state generation over realitic quantum networks. See more details in Sec. III.B of the Supplemental Material [59].
@@
It is known that $M=\sigma_{x}^{\otimes nd}$ is the optimal observable for $nd$-qubit pure GHZ probe state under unitary z-direction phase accumulation [60], and in principle each sensor node only needs to perform measurement of the local observable $\sigma_{x}^{\otimes n}$. However, this optimal measurement under ideal conditions is genuinely useless whenever there is any noise in the GHZ state as probe. Specifically, the variance of parameter estimation will diverge when taking the limit of small local parameters. We note that this conclusion is also implied by numerical results in another recent work by Cao and Wu [65]. In Sec. IV.A of the Supplemental Material [59] we analytically demonstrate this property for the depolarized GHZ state, while it holds generally for any GHZ-diagonal state with non-unit fidelity.
@@
Figure 2: Visualization of fidelity thresholds. The blue solid curve denotes the fidelity threshold given by the QFI, and the threshold given by optimized local azimuthal measurement is shown as the red solid curve. The horizontal and curved black dashed lines again represent the fidelity thresholds for $d$-qubit depolarized GHZ state to be GME and non-fully separable, respectively.
@@
The case of $n=1$ for Eq. 7 corresponds to noiseless local entanglement generation, similar to Eq. 5. We thus visualize and compare both fidelity thresholds in Fig. 2. The fidelity threshold in Eq. 7 is always higher than that given by Eq. 5, suggesting that the optimized local azimuthal measurement does not saturate the QCRB. On the other hand, although for small problem sizes $d<5$ the initial GHZ state needs to be GME to demonstrate quantum advantage if we use the optimized azimuthal measurement, when the problem size grows the requirement of initial fidelity drops, and in general the initial $d$-qubit GHZ state does not have to be GME. More details on the azimuthal measurement optimization can be found in Sec. IV.B of the Supplemental Material [59].
@@
Quantum network simulation.—Following the above analytics, we simulate GHZ state distribution given realistic first-generation quantum repeater [66, 67, 68] protocol stacks with an open-source, customizable quantum network simulator, SeQUeNCe [69]. We consider a simple 3-node network with linear topology, where a center node directly connects the other two end nodes through optical fibers, with a Bell state measurement station in the middle of each fiber link. First, bipartite entanglement links are established between the center node and the other nodes. Then LOCC, such as gate teleportation [70, 71, 72] and graph state fusion [73, 74], are performed to generate a GHZ state across all network nodes.
In our simulations we consider three scenarios, and thus three sets of system parameter values: $\{0.01\mathrm{s},0.1\mathrm{s},1\mathrm{s}\}$ for memory coherence times, $\{0.05,0.1,0.5\}$ for memory efficiency, and $\{0.8,0.85,0.9\}$ for raw entanglement fidelity. The first simulation scenario uses the first value in each of the above sets, and so on. Meanwhile, we fix other system parameters, especially: a 1s time interval allowed for entanglement distribution, a 10km distance between the center node and end nodes, 10 memories per end node and 20 memories on the center node so that entanglement purification [75, 76] is possible. We emphasize that the parameter values are not chosen from a specific reported experiment, but are selected according to the general vision of the state of the art and the potential future development of various candidate platforms for quantum networking, including solid state systems [28, 30, 77], atomic systems [29, 78], and superconducting systems [79]. Details of our quantum network simulation can be found in Sec. VII of the Supplemental Material [59].
@@
We have primarily focused on probe state preparation errors, while the parameter encoding dynamics can also be noisy due to sensor decoherence. It could be interesting to evaluate the impact of sensor decoherence after state preparation [81], where the time dependence of estimation accuracy gain [82, 83, 84] becomes important. We note that quantum error correction [85, 86, 87, 88] offers a good opportunity in fighting against noise in sensing process. Meanwhile, protocols such as continuous entanglement distribution [89, 90, 91] might be necessary to reduce latency of probe state preparation over quantum networks. In addition, optimization of Bell-state-based graph state distribution [92, 93, 94, 95, 96, 97, 98, 99, 100] is important for larger scale DQS. Notably, privacy and security [101, 102, 103] is a potentially important aspect of realistic quantum sensor network as well. Finally, although we mainly assume finite-dimensional matter-based quantum sensors, photonic systems [104, 105, 106, 107, 108, 109] are also playing an important role in DQS.

### v3

@@ Abstract
It is critically important to analyze the achievability of quantum advantage under realistic imperfections. In this work, we show that quantum advantage in distributed sensing can be achieved with noisy quantum networks which can only distribute noisy entangled states. We derive a closed-form expression of the quantum Fisher information (QFI) for estimating the average of local parameters using GHZ-diagonal probe states, an important distributed sensing prototype. From the QFI we obtain the necessary condition to achieve quantum advantage over the optimal local sensing strategy, which can also serve as an optimization-free entanglement detection criterion for multipartite states. In addition, we prove that genuine multipartite entanglement is neither necessary nor sufficient through explicit examples of depolarized and dephased GHZ states. We further explore the impacts from imperfect local entanglement generation and local measurement constraint, and our results imply that the quantum advantage is more robust against quantum network imperfections than local operation errors. Finally, we demonstrate that the probe state with potential for quantum advantage in distributed sensing can be prepared by a three-node quantum network using practical protocol stacks through simulations with SeQUeNCe, an open-source, customizable quantum network simulator.
@@
Introduction.—Distributed quantum sensing (DQS) [1, 2, 3] is one of the most important applications of quantum networks [4, 5]. It is expected to surpass classical sensing techniques in areas ranging from magnetometry [6, 7, 8, 9], phase imaging [10], precision clocks [11], energy applications [12], all the way to the exploration of fundamental physics [13, 14], including search for dark matter [15] and measuring stability of fundamental constants [16]. Over the past decade, DQS has attracted great theoretical interests [17, 18, 19, 20, 21, 22, 23, 24]. Various DQS protocols have been proposed under ideal conditions [25, 26, 27, 28, 29, 30, 31]. On the other hand, along with the experimental demonstration of concept in small-scale matter systems [32], the analysis of realistic DQS has also emerged recently, for instance, DQS with noise in signals [33, 34, 35, 36], atomic ensemble partition for DQS [37], and DQS with probabilistic quantum network operations [38]. However, the state-of-the-art progress in quantum networking [39, 40, 41] demonstrates that remote entanglement has much higher infidelity than local operations, so the process of initial state preparation will be the major error source for DQS in the near term. Despite its critical importance and necessity, the systematic analysis of DQS with imperfect state preparation surprisingly remains largely unexplored.
In this work, we focus on studying the possibility of realizing quantum advantage in DQS with noisy quantum networks which can only distribute noisy entangled states. We first specify the initial probe state, the parameter encoding process, and the estimation objective. We then derive the condition for quantum advantage over the optimal local sensing protocol, and demonstrate with the example of depolarized GHZ state as the initial probe. The impacts from imperfect local entanglement generation when scaling up the number of local sensors, and local measurement constraint are also analyzed. In addition, we perform a quantum network simulation which shows that DQS quantum advantage can in principle be achieved by realistic quantum network configurations.
@@
Under the aforementioned unitary encoding channel, the optimal probe state for this problem has been shown to be the global GHZ state  [2, 1], i.e., a GHZ state $|\mathrm{GHZ}\rangle=(|0\dots 0\rangle+|1\dots 1\rangle)/\sqrt{2}$ involving all the sensors on each sensor node. Therefore, we choose this global GHZ state as the target initial probe state to prepare in the quantum network of sensors. The preparation of the probe state over a quantum network can be decomposed into two steps. Firstly, the quantum network will distribute a $d$-qubit GHZ state across all $d$ sensor nodes, with each node having one qubit. Each node can then perform local entanglement generation [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53] to extend the size of the global GHZ state: By entangling $(n-1)$ additional quantum sensors per node, the global GHZ state will have $nd$ qubits in total. Note that only $(d-1)$ Bell pairs are needed to distribute the $d$-qubit GHZ state over the $d$ nodes (see Sec. VI of the Supplemental Material [54]). Therefore, such global GHZ states are easy to generate by quantum networks, which further justifies why we choose them as the probe states. Given noisy quantum networks, we consider that the prepared initial probe state is mixed. Motivated by Pauli twirling [55, 56, 57] which is able to transform error channels into Pauli channels, and stabilizer twirling [58] which can cancel off-diagonal density matrix elements under a stabilizer basis, we assume that the initial probe state is GHZ-diagonal: $\rho_{0}=\sum_{a}\lambda_{a}|\psi_{a0}\rangle\langle\psi_{a0}|$ with all $\lambda_{a}$ being non-zero, where $a$ is the index of the pure GHZ basis state $|\psi_{a0}\rangle$. This Pauli error model has also been utilized in a recent study of single-parameter quantum metrology with graph states [59]. According to the assumed GHZ-diagonal form of initial probe state, we can analytically derive the quantum Fisher information (QFI) [60] which characterizes the lower bound of parameter estimation variance, the quantum Cramér-Rao bound (QCRB).
Quantum Fisher information.—To derive the QFI for the average of all local parameters, we start with the QFI matrix (QFIM) [61, 62, 63, 64] for the $d$ local parameters. According to the assumed parametrization unitary, for a general GHZ-diagonal $nd$-qubit initial probe state $\rho_{0}$ where each of the $d$ sensor nodes holds $n$ qubit sensors, the QFIM for parameters $x$ is $\mathcal{F}(x)=(1-C)n^{2}J_{d}$, where $J_{d}$ is the $d$-by-$d$ matrix of ones, and $C$ captures the quality of the initial probe state, and it is analytically calculated as:
@@
Our parameter to estimate is $\theta_{1}=v_{1}^{T}x$ where $v\to(1,\dots,1)^{T}$. For concreteness, we follow the convention in [2, 1] and choose $v_{1}=(1,\dots,1)^{T}/\sqrt{d}$ which is normalized under the 2-norm. Then we are able to transform the QFIM for the new global parameter $\theta_{1}$, where the transformation could be operationally realized by constructing additional $(d-1)$ normalized vectors $v_{2},\dots,v_{d}$, s.t. $v_{i}^{T}v_{j}=\delta_{ij}$ [61]. Consequently, we have:
@@
where $\theta=(v_{1}^{T}x,\dots,v_{d}^{T}x)$, and the only non-zero entry is the QFI for $\theta_{1}$ of our interest: $\mathcal{F}(\theta_{1})=d(1-C)n^{2}$, which incorporates effects from both the network ($d$) and local resources ($n^{2}$). The details can be found in Sec. II of the Supplemental Material [54].
@@
It is well known that the GHZ state is the optimal probe state for local phase estimation [65], and indeed using separate GHZ states on each sensor node is the best local strategy (without any quantum communication between sensor nodes) for estimating the average of local parameters [2, 1]. The variance of this indirect estimation can be calculated through propagation of error [63]: $\mathrm{Var}_{\mathrm{local}}(\hat{\theta}_{1})=\sum_{l=1}^{d}\left(\frac{\partial\theta_{1}}{\partial x_{l}}\right)^{2}\mathrm{Var}_{\mathrm{local}}(\hat{x}_{l})$, where we use the QCRB of local GHZ state with $n$ qubits so that $\mathrm{Var}_{\mathrm{local}}(\hat{x}_{l})=1/(n^{2}\mu)$, where $\mu$ is the number of samples. Noticing that $\partial\theta_{1}/\partial x_{l}=1/\sqrt{d}$, we immediately have $\mathrm{Var}_{\mathrm{local}}(\hat{\theta}_{1})=1/(n^{2}\mu)$. Therefore, on top of the scaling advantage from increasing local resources, the additional relative quantum advantage of entangling sensor nodes with quantum networks is achievable when:
$\eta\equiv d(1-C)>1.$ (3)
Since entanglement in the initial probe state across sensor nodes is necessary for DQS quantum advantage, the above condition can also be interpreted as an operationally meaningful, optimization-free entanglement detection criterion for arbitrary $d$-qubit states. See details in Sec. II of the Supplemental Material [54]. Next, we will evaluate $\eta$ for $nd$-qubit GHZ state created from the initial noisy $d$-qubit GHZ state with noiseless and noisy local entanglement generation, respectively.
Noiseless local entanglement generation.—We first consider noiseless local entanglement generation, which can be interpreted as applying perfect CNOT gates between the qubit sensor initially entangled with other sensor nodes and other local qubit sensors prepared in $|0\rangle$ state to entangle, with the former being the control. This assumption results in that the $nd$-qubit GHZ state has the same fidelity as the $d$-qubit GHZ state. We are then allowed to decouple the network imperfections and local errors, and thus to understand the limit on the amount of network imperfections beyond which there cannot be any quantum advantage for DQS. While the analysis of $\eta$ is applicable to arbitrary GHZ-diagonal states, here for concreteness we assume a canonical full-rank error model for the initial $d$-qubit GHZ state, a depolarized GHZ state as a mixture of pure GHZ state and maximally mixed state, $\rho_{\mathrm{dp}}(F)$, where $F$ is the fidelity to the pure GHZ state. We can then substitute the $C$ for $d$-qubit depolarized GHZ state in Eq. 3 and let $\eta=1$ to derive a closed-form fidelity threshold, as a requirement on the quantum network performance for quantum advantage in DQS:
@@
We see that in the asymptotic regime of large $d$, the above fidelity threshold reduces to $1/d$. Moreover, the quantum advantage in DQS has a close relationship with quantum entanglement, and the entanglement properties of the depolarized GHZ state have been well studied: $d$-qubit depolarized GHZ states are not completely separable if $F>3/(2^{d}+2)$ [66], and genuine multipartite entangled (GME) if $F>1/2$ [67]. It can be shown that $F_{\mathrm{th,dp}}>3/(2^{d}+2)$ for $d\geq 2$, which means that entanglement is necessary for DQS quantum advantage. Meanwhile, except for the special case of $\left.F_{\mathrm{th,dp}}\right|_{d=3}\approx 0.50963>1/2$, $F_{\mathrm{th,dp}}<1/2$ for $d>3$, which suggests that GME is in principle unnecessary to demonstrate quantum advantage in DQS. Moreover, via analysis of rank-2 dephased GHZ states, we can show that GME is not sufficient for DQS quantum advantage either. Details can be found in Sec. III.A of the Supplemental Material [54].
@@
Figure 1: Visualization of $\eta_{\mathrm{dp}}$ as function of the number of local qubits $n$, for different initial fidelity $F$ and local entanglement generation quality $k$, with fixed $d=3$.
@@
Imperfect local entanglement generation.—We now start to include imperfections in local entanglement generation. Specifically, we consider the following phenomenological model to reflect the fidelity decrease when adding more qubits to the GHZ state: We assume that the state after local entanglement generation is an $nd$-qubit depolarized GHZ state, while the fidelity is modified according to the number of local qubits as $F(n)=k^{n-1}F$, where $F$ is the fidelity of the initial $d$-qubit GHZ state, and $k\in(0,1)$ is a parameter which describes the quality of local entanglement generation and thus the larger the better. The intuition behind this phenomenological model is the usage of noisy CNOT gates to generate GHZ states, and $k$ can be interpreted as gate fidelity. We can again evaluate $C$ in Eq. 1 as a function of the four system parameters $C_{\mathrm{dp}}(F,d,n,k)$, whose expression can be found in Sec. III.B of the Supplemental Material [54]. It can be shown that, as intuitively expected, $C_{\mathrm{dp}}$ decreases monotonically as initial fidelity $F$, number of local parameters $d$, and local entanglement generation quality $k$, increase for all $n\geq 1$, if $F,k>2^{-d}$. On the other hand, the dependence of $C_{\mathrm{dp}}$ on number of local quantum sensors $n$ is more complicated. We visualize $\eta_{\mathrm{dp}}=d(1-C_{\mathrm{dp}})$ with varying $n$ under different parameter choices in Fig. 1.
@@
From the plot we can verify that the relative advantage $\eta_{\mathrm{dp}}$ is larger when $F,k$ increase. We can also see non-monotonic behavior of $\eta_{\mathrm{dp}}$ when $k$ is sufficiently high in comparison to $F$, which indicates an increase in relative advantage for small local sensor numbers. The intersection point at which $\eta_{\mathrm{dp}}=1$ determines the maximal number of local quantum sensors $n_{\mathrm{max}}$ for quantum advantage over the optimal local sensing strategy to be potentially demonstrated. We observe that $n_{\mathrm{max}}$ does not change much when varying $F$ given fixed $k$, but it changes significantly when fixing $F$ and varying $k$, which suggests that the initial fidelity $F$ has less impact on $n_{\mathrm{max}}$ than the local entanglement generation quality $k$. This can be understood by considering the fidelity threshold for the $nd$-qubit depolarized GHZ state to be advantageous over the optimal local strategy, which also scales as $1/d$ and is almost independent of $n$. Imperfect local entanglement generation will decrease the GHZ fidelity to the threshold at $n_{\mathrm{max}}$. Therefore, we expect $k^{n_{\mathrm{max}}-1}F\approx 1/d$, which gives $n_{\mathrm{max}}\approx-\ln(dF)/\ln(k)$. This also implies robustness of the quantum advantage in DQS against the imperfection of probe state generation over realistic quantum networks. See more details in Sec. III.B of the Supplemental Material [54].
@@
It is known that $M=\sigma_{x}^{\otimes nd}$ is the optimal observable for $nd$-qubit pure GHZ probe state under unitary z-direction phase accumulation [65], and in principle each sensor node only needs to perform measurement of the local observable $\sigma_{x}^{\otimes n}$. However, if we use this measurement for noisy GHZ states, for instance depolarized GHZ state, the variance of parameter estimation diverges in the limit of small local parameters. We note that this conclusion is also implied by numerical results in another recent work by Cao and Wu [70]. In Sec. IV.A of the Supplemental Material [54] we analytically demonstrate this property for the depolarized GHZ state, while it holds generally for any GHZ-diagonal state with non-unit fidelity.
@@
Figure 2: Visualization of fidelity thresholds. The blue curve denotes the threshold given by the QFI, and the red curve denotes the threshold given by optimized local azimuthal measurement. The horizontal and curved black dashed lines represent the thresholds for $d$-qubit depolarized GHZ state to be GME and non-fully separable, respectively.
@@
The case of $n=1$ for Eq. 5 corresponds to noiseless local entanglement generation, similar to Eq. 4. We thus visualize and compare both fidelity thresholds in Fig. 2. The fidelity threshold in Eq. 5 is always higher than that given by Eq. 4, suggesting that the optimized local azimuthal measurement does not saturate the QCRB. On the other hand, although for small problem sizes $d<5$ the initial GHZ state needs to be GME to demonstrate quantum advantage if we use the optimized azimuthal measurement, when the problem size grows the requirement of initial fidelity drops, and in general the initial $d$-qubit GHZ state does not have to be GME. On the other hand, we can also show that the optimized local azimuthal measurement can saturate the QCRB for rank-2 dephased GHZ states, which implies an error model-dependent separability of optimal measurement for DQS. More details on the azimuthal measurement optimization can be found in Sec. IV.B of the Supplemental Material [54]. Moreover, we can use the GHZ state fidelity thresholds to estimate Bell state fidelity requirement, see Sec. IV.C of the Supplemental Material [54].
@@
Quantum network simulation.—Following the above analytics, we simulate GHZ state distribution given realistic first-generation quantum repeater [71, 72, 73, 74] protocol stacks with an open-source, customizable quantum network simulator, SeQUeNCe [75]. We consider a simple 3-node network with linear topology, where a center node directly connects the other two end nodes through optical fibers, with a Bell state measurement station in the middle of each fiber link. First, bipartite entanglement links are established between the center node and the other nodes. Then LOCC, such as gate teleportation [76, 77, 78] and graph state fusion [79, 80], are performed to generate a GHZ state across all network nodes.
In our simulations we consider three scenarios, and thus three sets of system parameter values: $\{0.01\mathrm{s},0.1\mathrm{s},1\mathrm{s}\}$ for memory coherence times, $\{0.05,0.1,0.5\}$ for memory efficiency, and $\{0.8,0.85,0.9\}$ for raw entanglement fidelity. The first simulation scenario uses the first value in each of the above sets, and so on. Meanwhile, we fix other system parameters, especially: a 1s time interval allowed for entanglement distribution, a 10km distance between the center node and end nodes, 10 memories per end node and 20 memories on the center node so that entanglement purification [81, 82] is possible. We emphasize that the parameter values are not chosen from a specific reported experiment, but are selected according to the general vision of the state of the art and the potential future development of various candidate platforms for quantum networking, including solid state systems [39, 41, 83], atomic systems [40, 84], and superconducting systems [85]. The network topology is visualized in Fig. 3. Details of our quantum network simulation can be found in Sec. VII of the Supplemental Material [54].
Figure 3: Topology of the simulated 3-node quantum network.
@@
We have primarily focused on probe state preparation errors because it is the major error source in quantum networks. It could also be interesting to evaluate the impact of sensor decoherence after state preparation [87], where the time dependence of estimation accuracy gain [88, 89, 90] becomes important. We note that quantum error correction [91, 92, 93, 94] offers a good opportunity in fighting against noise in sensing process. Meanwhile, protocols such as continuous entanglement distribution [95, 96, 97, 98, 99] might be necessary to reduce latency of probe state preparation over quantum networks. In addition, optimization of Bell-state-based graph state distribution [100, 101, 102, 103, 104, 105, 106, 107, 108] is important for larger scale DQS. Notably, privacy and security [109, 110, 111] is a potentially important aspect of realistic quantum sensor network as well. Finally, although we mainly assume finite-dimensional matter-based quantum sensors, photonic systems [112, 113, 114, 115, 116, 117] are also playing an important role in DQS.